# Incentive-Aligned Multi-Source LLM Summaries

### v1

@@ Methods.
In all our experiments, LLM judges are run using gemini-2.5-flash (Comanici et al., 2025) to make the results comparable. In the main experimental section, we presented experiment where the source generation uses gemini-2.5-flash and the claim extraction pipeline uses gemini-2.5-flash-lite (Comanici et al., 2025).
We justified that even though the mechanism itself is model-agnostic, we chose, in the main text, the Flash model for generation of the synthetic sources, and the lightweight model to prioritize the low latency and efficiency required for search applications, reflecting a realistic asymmetry where attackers can expend more effort than a real-time defense. We now expand the analysis by expanding to two additional variants, (1) source generation uses gemini-2.5-flash and the claim extraction pipeline uses gemini-2.5-flash, (2) source generation uses gemini-2.5-flash-lite and the claim extraction pipeline uses gemini-2.5-flash-lite. We aim to show that beyond the setup of the results presented in the main text, and across different models, our method achieves significant improvement over baselines.
I.8.1 Results 1: Robustness against adversarial and untruthful sources
@@
On the ClashEval dataset, the performance gap between our method and the baselines is even more stark. Baseline methods struggle significantly, with answer accuracies often in the single or low double digits. In contrast, TTS consistently achieves over 68
@@
I.8.2 Results 2: Robustness against coordinated, uninformative behavior

### v2

@@ Methods.
In all our experiments, LLM judges are run using gemini-2.5-flash [Comanici et al., 2025] to make the results comparable. In the main experimental section, we presented experiment where the source generation uses gemini-2.5-flash and the claim extraction pipeline uses gemini-2.5-flash-lite [Comanici et al., 2025].
We now expand the analysis by expanding to two additional variants, (1) source generation uses gemini-2.5-flash and the claim extraction pipeline uses gemini-2.5-flash, (2) source generation uses gemini-2.5-flash-lite and the claim extraction pipeline uses gemini-2.5-flash-lite. We justify that we chose the lightweight model to prioritize the low latency and efficiency required for search applications, though the mechanism itself is model-agnostic. This also reflects a realistic asymmetry where attackers can expend more effort than a real-time defense. Here, we aim to show that even without this asymmetry, and across different models, our method achieve significant improvement over baselines.
H.9.1 Results 1: Robustness against adversarial and untruthful sources
@@
On the ClashEval dataset, the performance gap between our method and the baselines is even more stark. Baseline methods struggle significantly, with answer accuracies often in the single or low double digits. In contrast, TTS consistently achieves over 68% answer accuracy, demonstrating its effectiveness in handling the challenging, multi-domain nature of this benchmark.
@@
H.9.2 Results 2: Robustness against coordinated, uninformative behavior